プリンキピア

第7章中心極限定理

Exercises 91–97練習問題 91–97

For each problem, wherever possible, provide graphs and use the calculator.
各問題では、可能な限りグラフを示し、計算機を使う。
Exercise 91問題 91
NeverReady batteries has engineered a newer, longer lasting AAA battery. The company claims this battery has an average life span of 17 hours with a standard deviation of 0.8 hours. Your statistics class questions this claim. As a class, you randomly select 30 batteries and find that the sample mean life span is 16.7 hours. If the process is working properly, what is the probability of getting a random sample of 30 batteries in which the sample mean lifetime is 16.7 hours or less? Is the company’s claim reasonable?
英語のヒント
ネヴァーレディ電池社は、より長持ちする新しい単4電池を開発した。会社は、この電池の平均寿命は17時間、標準偏差は0.8時間だと主張している。あなたの統計学のクラスは、この主張に疑問を持つ。クラスで電池30本を無作為に選ぶと、標本平均寿命は16.7時間であった。製造工程が適切に機能しているなら、電池30本の無作為標本で、標本平均寿命が16.7時間以下となる確率はいくつか。会社の主張は妥当か。
Solution解き方
  • We have μ = 17, σ = 0.8, x ¯ = 16.7, and n = 30. To calculate the probability, we use normalcdf(lower, upper, μ, σ n ) = normalcdf ( E–99,16.7,17, 0.8 30 ) = 0.0200.
    μ = 17、σ = 0.8、 x ¯ = 16.7、n = 30である。確率を求めるには、normalcdf(下限, 上限, μ, σ n )= normalcdf ( E–99,16.7,17, 0.8 30 ) = 0.0200を使う。
  • If the process is working properly, then the probability that a sample of 30 batteries would have at most 16.7 lifetime hours is only 2%. Therefore, the class was justified to question the claim.
    製造工程が適切に機能しているなら、30本の電池の標本の平均寿命が16.7時間以下となる確率は、わずか2%である。したがって、このクラスが会社の主張を疑ったのには正当な理由があった。
Exercise 92問題 92
Men have an average weight of 172 pounds with a standard deviation of 29 pounds.
男性の平均体重は172ポンド、標準偏差は29ポンドである。
  1. Find the probability that 20 randomly selected men will have a sum weight greater than 3600 lbs.
    無作為に選んだ男性20人の体重の合計が3,600ポンドを超える確率を求める。
  2. If 20 men have a sum weight greater than 3500 lbs, then their total weight exceeds the safety limits for water taxis. Based on (a), is this a safety concern? Explain.
    男性20人の体重の合計が3,500ポンドを超えると、水上タクシーの安全上の制限を超過する。aの結果に基づくと、これは安全上の懸念となるか。説明する。
Exercise 93問題 93
M&M candies large candy bags have a claimed net weight of 396.9 g. The standard deviation for the weight of the individual candies is 0.017 g. The following table is from a stats experiment conducted by a statistics class.
英語のヒント
M&Mチョコレートの大袋には、正味重量396.9 gと表示されている。個々のチョコレートの重量の標準偏差は0.017 gである。次の表は、統計学のクラスで行った統計実験の結果である。
Red
赤
Orange
オレンジ
Yellow
黄
Brown
茶
Blue
青
Green
緑
0.751
0.735
0.883
0.696
0.881
0.925
0.841
0.895
0.769
0.876
0.863
0.914
0.856
0.865
0.859
0.855
0.775
0.881
0.799
0.864
0.784
0.806
0.854
0.865
0.966
0.852
0.824
0.840
0.810
0.865
0.859
0.866
0.858
0.868
0.858
1.015
0.857
0.859
0.848
0.859
0.818
0.876
0.942
0.838
0.851
0.982
0.868
0.809
0.873
0.863
0.803
0.865
0.809
0.888
0.932
0.848
0.890
0.925
0.842
0.940
0.878
0.793
0.832
0.833
0.905
0.977
0.807
0.845
0.850
0.841
0.852
0.830
0.932
0.778
0.856
0.833
0.814
0.842
0.881
0.791
0.778
0.818
0.810
0.786
0.864
0.881
0.853
0.825
0.864
0.855
0.873
0.942
0.880
0.825
0.882
0.869
0.931
0.912
0.887
Table 7.1表 7.1
The bag contained 465 candies and he listed weights in the table came from randomly selected candies. Count the weights.
袋には465粒が入っており、表の重量は無作為に選んだ粒のものである。重量データの個数を数える。
  1. Find the mean sample weight and the standard deviation of the sample weights of candies in the table.
    表のチョコレートの標本平均重量と、標本の重量の標準偏差を求める。
  2. Find the sum of the sample weights in the table and the standard deviation of the sum of the weights.
    表の標本の重量の和と、重量の和の標準偏差を求める。
  3. If 465 M&Ms are randomly selected, find the probability that their weights sum to at least 396.9.
    M&Mを465粒無作為に選んだ場合、その重量の和が少なくとも396.9となる確率を求める。
  4. Is the Mars Company’s M&M labeling accurate?
    マース社のM&Mの表示は正確か。
Solution解き方
  1. For the sample, we have n = 100, x ¯ = 0.862, s = 0.05
    この標本では、n = 100、 x ¯ = 0.862、s = 0.05である。
  2. Σ x ¯ = 85.65, Σs = 5.18
  3. normalcdf(396.9,E99,(465)(0.8565),(0.05)( 465 )) ≈ 1
  4. Since the probability of a sample of size 465 having at least a mean sum of 396.9 is appproximately 1, we can conclude that Mars is correctly labeling their M&M packages.
    大きさ465の標本で平均的な合計重量が少なくとも396.9となる確率は、ほぼ1なので、マース社はM&Mの包装に正しく表示していると結論できる。
Exercise 94問題 94
The Screw Right Company claims their 3 4 inch screws are within ±0.23 of the claimed mean diameter of 0.750 inches with a standard deviation of 0.115 inches. The following data were recorded.
英語のヒント
スクリュー・ライト社は、同社の 3 4 インチのねじについて、直径は表示上の平均0.750インチの±0.23以内、標準偏差は0.115インチであると主張している。次のデータが記録された。
0.757
0.723
0.754
0.737
0.757
0.741
0.722
0.741
0.743
0.742
0.740
0.758
0.724
0.739
0.736
0.735
0.760
0.750
0.759
0.754
0.744
0.758
0.765
0.756
0.738
0.742
0.758
0.757
0.724
0.757
0.744
0.738
0.763
0.756
0.760
0.768
0.761
0.742
0.734
0.754
0.758
0.735
0.740
0.743
0.737
0.737
0.725
0.761
0.758
0.756
Table 7.2表 7.2
The screws were randomly selected from the local home repair store.
ねじは地元の住宅修理用品店から無作為に選んだ。
  1. Find the mean diameter and standard deviation for the sample
    標本の平均直径と標準偏差を求める。
  2. Find the probability that 50 randomly selected screws will be within the stated tolerance levels. Is the company’s diameter claim plausible?
    英語のヒント
    無作為に選んだ50本のねじが、示された許容範囲に入る確率を求める。会社が主張する直径はもっともらしいか。
Exercise 95問題 95
Your company has a contract to perform preventive maintenance on thousands of air-conditioners in a large city. Based on service records from previous years, the time that a technician spends servicing a unit averages one hour with a standard deviation of one hour. In the coming week, your company will service a simple random sample of 70 units in the city. You plan to budget an average of 1.1 hours per technician to complete the work. Will this be enough time?
英語のヒント
あなたの会社は、大都市にある数千台のエアコンの予防保守を行う契約を結んでいる。過去の保守記録によると、技術者が1台の整備にかける時間は平均1時間、標準偏差1時間である。来週、会社は市内の70台からなる単純無作為標本を整備する予定である。仕事を終えるために、技術者1人あたり平均1.1時間を見込んでいる。これで十分な時間になるか。
Solution解き方
Use normalcdf ( E–99,1.1,1, 1 70 ) = 0.7986. This means that there is an 80% chance that the service time will be less than 1.1 hours. It could be wise to schedule more time since there is an associated 20% chance that the maintenance time will be greater than 1.1 hours.
normalcdf ( E–99,1.1,1, 1 70 ) = 0.7986を使う。これは、保守作業の時間が1.1時間未満となる確率が80%であることを意味する。保守時間が1.1時間を超える確率も20%あるので、もっと時間を見込むほうが賢明かもしれない。
Exercise 96問題 96
A typical adult has an average IQ score of 105 with a standard deviation of 20. If 20 randomly selected adults are given an IQ test, what is the probability that the sample mean scores will be between 85 and 125 points?
一般的な成人のIQ得点の平均は105、標準偏差は20である。無作為に選んだ成人20人にIQ検査を行うとき、標本平均得点が85点から125点の間にある確率はいくつか。
Exercise 97問題 97
Certain coins have an average weight of 5.201 grams with a standard deviation of 0.065 g. If a vending machine is designed to accept coins whose weights range from 5.111 g to 5.291 g, what is the expected number of rejected coins when 280 randomly selected coins are inserted into the machine?
英語のヒント
ある硬貨の平均重量は5.201 g、標準偏差は0.065 gである。重量が5.111 gから5.291 gの硬貨を受け付けるように自動販売機が設計されているとする。無作為に選んだ硬貨280枚を投入したとき、返却される硬貨の枚数の期待値はいくつか。
Solution解き方
We assume that the weights of coins are normally distributed in the population. Since we have normalcdf ( 5.111,5.291,5.201, 0.065 280 ) ≈ 0.8338, we expect (1 – 0.8338)280 ≈ 47 coins to be rejected.
母集団の硬貨重量は正規分布に従うと仮定する。normalcdf ( 5.111,5.291,5.201, 0.065 280 ) ≈ 0.8338なので、返却される枚数の期待値は(1 − 0.8338)280 ≈ 47枚となる。
Exponential Distribution
指数分布
a continuous random variable (RV) that appears when we are interested in the intervals of time between some random events, for example, the length of time between emergency arrivals at a hospital, notation: X ~ Exp(m). The mean is μ = 1 m and the standard deviation is σ = 1 m . The probability density function is f(x) = me–mx, x ≥ 0 and the cumulative distribution function is P(X ≤ x) = 1 – e–mx.
病院への救急患者の到着間隔など、無作為な事象の間の時間間隔に関心があるときに現れる連続確率変数。表記はX ~ Exp(m)。平均はμ = 1 m 、標準偏差はσ = 1 m である。確率密度関数はf(x) = me–mx、x ≥ 0であり、累積分布関数はP(X ≤ x) = 1 − e–mxである。
Mean
平均
a number that measures the central tendency; a common name for mean is "average." The term "mean" is a shortened form of "arithmetic mean." By definition, the mean for a sample (denoted by x ¯ ) is x ¯ = Sum of all values in the sample Number of values in the sample , and the mean for a population (denoted by μ) is μ = Sum of all values in the population Number of values in the population .
英語のヒント
中心的な傾向を測る数値。meanは一般にaverageとも呼ばれる。meanはarithmetic mean(算術平均)の略である。定義により、 x ¯ で表す標本の平均は x ¯ = 標本内の全数値の和 標本内の数値の個数 、μで表す母集団の平均は μ = 母集団内の全数値の和 母集団内の数値の個数 である。
Normal Distribution
正規分布
a continuous random variable (RV) with pdf f(x) = 1 σ 2π e – (x – μ) 2 2 σ 2 , where μ is the mean of the distribution and σ is the standard deviation.; notation: X ~ N(μ, σ). If μ = 0 and σ = 1, the RV is called the standard normal distribution.
確率密度関数が f(x) = 1 σ 2π e – (x – μ) 2 2 σ 2 である連続確率変数。ここでμは分布の平均、σは標準偏差である。表記はX ~ N(μ, σ)。μ = 0、σ = 1の場合、その確率変数の分布を標準正規分布という。
Uniform Distribution
一様分布
a continuous random variable (RV) that has equally likely outcomes over the domain, a < x < b; often referred as the Rectangular Distribution because the graph of the pdf has the form of a rectangle. Notation: X ~ U(a, b). The mean is μ = a + b 2 and the standard deviation is σ = (b–a) 2 12 . The probability density function is f(x) = 1 b–a for a < x < b or a ≤ x ≤ b. The cumulative distribution is P(X ≤ x) = x–a b–a .
英語のヒント
定義域a < x < bにおいて、結果が等しく起こりうる連続確率変数。確率密度関数のグラフが長方形になるため、矩形分布とも呼ばれる。表記はX ~ U(a, b)。平均は μ = a + b 2 、標準偏差は σ = (b–a) 2 12 である。確率密度関数は、a < x < bまたはa ≤ x ≤ bに対して f(x) = 1 b–a である。累積分布はP(X ≤ x) = x–a b–a である。